OpenStudy (anonymous):

Solve using matrices x+z=7 2y-z=-3 -x-3y+2z=11

OpenStudy (unklerhaukus):

\[\left[\begin{array}{ccc}1&0&1\\0&2&-1\\-1&-3&2\end{array}\right|\left.\begin{array}{ccc}7\\-3\\11\end{array}\right]\]